OpenStudy (anonymous):

Solve. Write the answer in simplest form. (Must show work!) 1. 1/2x = 5/8

OpenStudy (anonymous):

\[\frac{1}{2}x=\frac{5}{8}\] or \[\frac{1}{2x}=\frac{5}{8}\]??

OpenStudy (anonymous):